• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Ramesh

    Log entry time 08:05:36 on March21,2005

    Entry number 142079

    keyword=Shift Summary

    Shift Crew: Eli Piasetzky, Ramesh Subedi.


    HRS angles: left 19.5, right 40.
    HRS momentum settings: left 3.724 MeV/c, right 1.45 MeV/c
    BigBite/Neutron Array angle: -99


    00:00 Shift started with the run 2836 from swing shift.

    All runs Production Runs in C-tilted at 40uA beam.

    00:30-01:14 Run 2837(200k)
    01:14-01:55 Run 2838(200k)
    01:55-02:17 Run 2839(99k)
    02:17 TS2 (Trigger Supervisor of SRC) failed, ended the run 2839 prematurely (run-end failed). Rebooting the TS2 worked fine.

    02:24-03:06 Run 2840(200k)
    03:06-03:49 Run 2841(200k)
    03:50-04:19 Run 2842(110k) stopped early due to beam off for a while.
    04:19 MCC told that beam will be off for a while due to RF-problem.
    04:59 beam back

    05:00-05:43 Run 2843(200k)
    05:44-06:29 Run 2844(200k)
    06:30-07:13 Run 2845(200k)
    07:14 Run 2846 passed to day shift.